OpenStudy (anonymous):

@iGreen 1 more after this one and then we will probably be done. Thanks again

OpenStudy (anonymous):

A ball is made of a plastic sphere covered with a layer of rubber. The plastic sphere has a radius of 18 in., and the ball itself has a radius of 21 in. What is the volume of the rubber layer? Use 3.14 to approximate pi and express your final answer in hundredths. ____ in^3

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

If the radius of the whole thing is 21, and the plastic part is 18, subtract 21 - 18 to get the radius of the rubber layer.

OpenStudy (anonymous):

21 - 18 = 3

OpenStudy (igreen):

Find the two volumes, subtract.

OpenStudy (igreen):

Or we could do it that way.

OpenStudy (anonymous):

>:3

OpenStudy (igreen):

\(V = \dfrac{4}{3} \pi r^3\) Plug in 3 for the radius

OpenStudy (anonymous):

3^3 = 27

OpenStudy (igreen):

Yep, now multiply by 4/3 and pi

OpenStudy (anonymous):

113.04

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

1 more after this one and odds are we will be done. Thanks again

OpenStudy (igreen):

No problem.